# Provably Unlearnable Examples

### v1

@@ 3. Problem Statement
Definition 3.1 (True Learnability of a perturbed dataset).
@@
In essence, the problem to be addressed in this paper is two-fold. First, we certify the learnability of a dataset towards arbitrary classifiers trained on it as long as the trained parameters are from $\hat{\Theta}$. This setting can appear in cases such as transfer learning, model fine-tuning, and recovery attacks. Second, we strive to reduce the certified learnability of the protected dataset by designing a proper $\delta$. We will present in the following part a threat model of the paper.
Threat model.  The producer of UEs (i.e., defender) has white-box access to a surrogate classifier $f_{\hat{\theta}}$ and a public dataset ${\mathbb{D}}_{s}\sim{\mathcal{D}}$. Moreover, either a to-be-protected private clean test dataset ${\mathbb{D}}$ or the domain ${\mathcal{D}}$ is revealed to the defender. The defender can alter the parameters of $f_{\hat{\theta}}$ and modify ${\mathbb{D}}_{s}$ at will. The defender crafts an unlearnable version ${\mathbb{D}}_{s}\oplus\delta$ of ${\mathbb{D}}_{s}$ and releases the unlearnable dataset rather than the unprotected one to the public. Besides, defenders can either train surrogate classifiers on generated ${\mathbb{D}}_{s}\oplus\delta$ or simultaneously with $\delta$ on ${\mathbb{D}}_{s}$. However, the defender has zero access to any potential unauthorized classifiers and is unaware of the training procedure for the unauthorized classifiers.
@@
Among these adversaries, GA encompasses almost all attackers in real-world scenarios. CA denotes the subset of attackers from GA that can be provably mitigated by our defense. SA serves as a tool to gauge the robustness of different UEs against GAs who cannot be provably invalidated, as well as to assess the tightness of our defense against CAs.
4. Overview
We will provide a brief overview over the crux of learnability certification and PUE generation in this section. Our first desideratum is measuring how well an unlearnable dataset can sabotage those illegal but unknown classifiers by certifying a possibly best learning outcome on this dataset. In a nutshell, a surrogate classifier selected based on the unlearnable set ${\mathbb{D}}_{s}\oplus\delta$ can be viewed as a hypothesis of the data distribution. Intuitively, if the distribution of the unlearnable set is highly disjoint from that of clean data, the classifier would demonstrate a significant generalization error on the clean test dataset. In contrast, if the unlearnable set is suboptimal, the classifier might still be able to generalize to the mapping relationships in the clean dataset and thus induce a lower generalization error. Based on the aforementioned intuition, we convert the certification problem to certifying a surrogate classifier appropriately modeling the distribution of UEs.
Second, we discuss and design a simple method for selecting the surrogate classifier used in the certification. Importantly, we emphasize that the two entities, the UE and the surrogate classifier, serve different objectives. On the side of UEs, their purpose is upfront — to minimize the learnability by lowering the clean testing accuracy of any classifiers trained on them. However, the design goal of the surrogate classifier diverges from that of UEs. Instead, the surrogate should 1) resemble the distributional characteristics of UEs and 2) have a high chance of discovering better clean test accuracy when its parameters are perturbed. In lieu of directly using classifiers trained by UEs as surrogates, we design a better way to construct surrogates meeting these two criteria.
@@ 5. Learnability Certification
5.1. Quantile Parametric Smoothing
@@
Definition 5.1 (Quantile Parametric Smoothing function).
@@
(7) $h_{q}(\theta)=\inf\ \{t\ |\ \Pr_{\epsilon}[A_{{\mathbb{D}}}(\theta+\epsilon)\leq t]\geq q\},$
@@
The definition connects to percentile smoothing used in certified robustness (Chiang et al., 2020; Bansal et al., 2022). In lieu of randomizing the inputs of classifiers, QPS takes the parameters (i.e., weights) of classifiers as its variables and randomizes the parameters.
@@
5.2. Certified Learnability
To apply the QPS function to obtain a learnability certification, we explicitly define $A_{{\mathbb{D}}}$ as a performance metric in this section. Without loss of generality, such $A_{{\mathbb{D}}}(\theta)$ can be the top-1 accuracy function of a classifier ${f}_{\hat{\theta}}$ mapping its input $x$ to a label $y$, where $(x,y)\in{\mathbb{D}}$. That is
(8) $A_{{\mathbb{D}}}(\hat{\theta}):=\mathbb{E}_{(x,y)\sim{\mathbb{D}}}\mathds{1}[{f}_{\hat{\theta}}(x)=y].$
@@ 5.2. Certified Learnability
Figure 3. An illustration of the certified $(q,\eta)$-Learnability.
@@
Theorem 5.2 (Perturbation bound on QPS).
@@
(9) ${h}_{q}(\hat{\theta}+\upsilon)\leq\inf\ \{t\ |\ \Pr_{\epsilon}[A_{{\mathbb{D}}}(\hat{\theta}+\epsilon)\leq t]\geq\overline{q}\},\ \forall\ \|\upsilon\|\leq\eta,$
@@
We defer the proof of Theorem 5.2 to Appendix B. The theorem states that the return of the QPS function is bounded with respect to the perturbation $\upsilon$ in its input. Recall that ${h}_{q}(\hat{\theta}+\upsilon)$ is actually an upper bound of the clean test accuracy for classifiers whose parameters $\theta\sim\mathcal{N}(\hat{\theta}+\upsilon,\sigma^{2}I)$, with probability at least $q$. Therefore, the right-hand side of Inequality 9 actually guarantees the best possible clean test accuracy for classifiers with parameters from the subspace $\hat{\Theta}:=\{\theta\ |\ \theta\sim\mathcal{N}(\hat{\theta}+\upsilon,\sigma^{2}I),\ \|\upsilon\|\leq\eta\}$. We name such $\hat{\Theta}$ as a certified parameter set. Though it seems that $\hat{\Theta}$ covers all parameters in $\mathbb{R}^{d}$, we only consider those $\theta$ having a sufficiently large probability of being sampled.
Based on Theorem 5.2, we have the formal definition of $(q,\eta)$-Learnability as follows.
Definition 5.3 ($(q,\eta)$-Learnability).
@@
(10) $l_{(q,\eta)}(\hat{\Theta};{\mathbb{D}}_{s}\oplus\delta)=\inf\ \{t|\Pr_{\epsilon}[A_{{\mathbb{D}}}(\hat{\theta}+\epsilon)\leq t]\geq\overline{q}\},$
@@
It should be noted that, though ${\mathbb{D}}_{s}\oplus\delta$ does not explicitly appear in the function, $\hat{\theta}$ is selected by a learning function $\hat{\theta}=\Gamma({\mathbb{D}}_{s}\oplus\delta)$ given ${\mathbb{D}}_{s}\oplus\delta$. Henceforth, $\hat{\theta}$ can be viewed as a function of ${\mathbb{D}}_{s}\oplus\delta$ in this case. In practice, we can sample $A_{{\mathbb{D}}}(\hat{\theta}+\epsilon)$ using Monte Carlo and obtain the accuracy in the empirical $\overline{q}$-th quantile. Next, a Binomial confidence interval upper bound of the accuracy can be adopted as $l_{(q,\eta)}(\hat{\Theta};{\mathbb{D}}_{s}\oplus\delta)$ with a confidence level of $1-\alpha$. Detailed steps of computing the confidence interval upper bound are in Appendix A.
@@
5.3. Certification Algorithm
We summarize the certification algorithm in this section. We first train a surrogate classifier $\hat{f}_{\hat{\theta}}$ of which the parameters $\hat{\theta}$ will be randomized for $n$ times over a Gaussian distribution to calculate the QPS function. For each randomization, the resulting classifier with parameters $\hat{\theta}+\epsilon$ is evaluated on the clean test set ${\mathbb{D}}$ such that $A_{{\mathbb{D}}}(\hat{\theta}+\epsilon)$ is obtained. Next, $\overline{q}$ can be theoretically computed based on Theorem 5.2. Finally, we compute the confidence interval of the $A_{{\mathbb{D}}}(\hat{\theta}+\epsilon)$ in the $\overline{q}$-th quantile to obtain the upper bound of the interval as $l_{(q,\eta)}(\hat{\Theta};{\mathbb{D}}_{s}\oplus\delta)$. The detailed certification process is depicted in Algorithm 1 and Algorithm 2. In addition, a meticulous description is in Appendix A.
There are a few important things to be noticed. First, the $\sigma$ of the parametric smoothing noise should be selected with caution. There is an intuition behind randomizing the surrogate weights and computing the QPS function — some of the randomized classifiers can outperform the surrogate classifier on the clean test set. Notwithstanding, we find that, when an improperly large $\sigma$ is employed, the number of randomized classifiers having a higher accuracy than that of the surrogate drops significantly. The possible reason for this phenomenon could be that the shortcuts added to the UEs cannot fully conceal the mapping information between input samples and their labels. Hence, the surrogate trained on the unlearnable dataset, though not adequately, is somehow fit to the clean data distribution as well. Large noises added to the parameters thus not only shift the surrogate away from the distribution of UEs but also move it off the clean data manifold. This assumption is further verified by the recovery attack, which discovers classifiers with satisfactory performance on clean data in the vicinity of the poisoned classifier. Therefore, we further introduce the methods for constructing sound surrogates for tighter certification in Section 6.
1 func QUpperBound
@@ 5.3. Certification Algorithm
2 $\overline{q}\leftarrow\Phi(\Phi^{-1}(q)+\frac{\eta}{\sigma})$
3 $\underline{k},\ \overline{k}\ \leftarrow\ \lceil n*\overline{q}\rceil,\ n$
@@
5 for $k\in\{\underline{k},\underline{k}+1,...,\overline{k}\}$ do
6       if $\textsc{Binomial}(n,k,\overline{q})>1-\alpha$ then
7             $k^{*}\ \leftarrow k$
8
9      else
10             $Continue$
11
12
13if $k^{*}\neq 0$ then
@@ 10             $Continue$
14
15else
16       Abstain
17func Binomial
@@ 16       Abstain
18 $\textsc{Conf}\ \leftarrow\ \sum_{i=1}^{k}\binom{n}{k}(\overline{q})^{i}(1-\overline{q})^{n-i}$
@@
1 Initialize $a$
2 for $i\in 1,...,n$ do
3       $\theta\ \leftarrow\ \hat{\theta}+\epsilon$, $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$
4       Evaluate $A_{{\mathbb{D}}}({\theta})$ on ${\mathbb{D}}$
5       Append $A_{{\mathbb{D}}}({\theta})$ to $a$
6
[…]
8 $k\ \leftarrow\ \textsc{QUpperBound}(n,\alpha,\sigma,\eta,q)$
9 $t\ \leftarrow\ a_{k}$
@@ 5.4. Generalization of the Certification
As a solution, the defender can sample a test set $\hat{{\mathbb{D}}}$ from the accessible domain ${\mathcal{D}}$ and certify a generalization $(q,\eta)$-Learnability which bounds the learnability on test sets drawn from ${\mathcal{D}}$. By Hoeffding’s Inequality, $A_{{\mathcal{D}}}(\hat{\theta})\leq A_{\hat{{\mathbb{D}}}}(\hat{\theta})+\sqrt{\frac{1}{2N}\log(\frac{2}{\beta})}$ exists with probability at least $1-\beta$. The right-hand side of the inequality can be used in the QPS function instead of $A_{{\mathbb{D}}}(\hat{\theta})$. Combining with the union bound, a certified generalization $(q,\eta)$-Learnability can be acquired as $l_{(q,\eta)}(\hat{\Theta};{\mathbb{D}}_{s}\oplus\delta)+\sqrt{\frac{1}{2N}\log(\frac{2n}{\beta})}$, where $n$ is the number of noise draws. Moreover, note that the classifiers sampled around a surrogate relate to the PAC-Bayesian predictors (e.g., Gibbs classifiers) whose parameters are randomly drawn from a distribution (McAllester, 2003). We can thus bound the expected generalization accuracy under random weight perturbation through the following corollary.
Corollary 5.4 (Expected generalization accuracy under parametric smoothing noise).
@@
(12) $\mathbb{E}_{\epsilon}[A_{{\mathcal{D}}}(\hat{\theta}+\epsilon)]\geq\mathbb{E}_{\epsilon}[{A}_{\hat{{\mathbb{D}}}}(\hat{\theta}+\epsilon)]-\sqrt{\frac{\frac{\|\hat{\theta}\|^{2}}{\sigma^{2}}+\ln\frac{N}{\alpha}}{2(N-1)}},$
@@
The corollary indicates that the expected generalization accuracy over the clean data distribution is lower bounded by the expected empirical accuracy minus a function of $\hat{\theta}$, $\sigma$ and $N$. Despite that the corollary bounds the expectation instead of the $q$-th quantile, it connotes that a $f_{\hat{\theta}}$ constantly producing correct predictions on $\hat{{\mathbb{D}}}$ also has a high generalization accuracy, which evinces the inferiority of the UEs resulting $\hat{\theta}$. Furthermore, if the expectation of ${A}_{\hat{{\mathbb{D}}}}(\hat{\theta}+\epsilon)$ is robust to varying $\sigma$, the expected generalization accuracy would grow with increasing $\sigma$. For appropriately skewed accuracy distributions, this result implies that an $f_{\hat{\theta}}$ which endures large parametric noise can certify $(q,\eta)$-Learnability that generalizes better to ${\mathcal{D}}$.
@@
[backgroundcolor=grey!10,rightline=true,leftline=true,topline=true,bottomline=true,roundcorner=1mm,everyline=false,nobreak=false] Remark.  We defined the QPS function of accuracy scores obtained by a set of classifiers randomized from a surrogate classifier fit to UEs ${\mathbb{D}}_{s}\oplus\delta$. The QPS function returns the value in the $q$-th quantile of all accuracy scores. The output of the QPS function can be further upper bounded by Theorem 5.2. The upper bound of the accuracy in the $\overline{q}$-th quantile is treated as the $(q,\eta)$-Learnability of ${\mathbb{D}}_{s}\oplus\delta$. The key message of the certified $(q,\eta)$-Learnability is that — if an unauthorized classifier has its trained parameters in the certified parameter set $\hat{\Theta}:=\{\theta\ |\ \theta\sim\mathcal{N}(\hat{\theta}+\upsilon,\sigma^{2}I),\ \|\upsilon\|\leq\eta\}$, then its clean test accuracy will not exceed the $(q,\eta)$-Learnability, with probability at least $q$. Moreover, if ${\mathbb{D}}$ is hidden from the defender, a generalization $(q,\eta)$-Learnability can be certified by sampling a test set from ${\mathcal{D}}$ by the defender.
6. Provably Unlearnable Examples

### v2

@@ III Problem Statement
Definition 1 (True Learnability of a perturbed dataset).
@@
In essence, the problem to be addressed in this paper is two-fold. First, we certify the learnability of a dataset towards arbitrary classifiers trained on it as long as the trained parameters are from $\hat{\Theta}$. This setting non-trivially appears in data exploitation cases through model retraining, transfer learning, model fine-tuning, and recovery attacks. Second, we strive to reduce the learnability of the protected dataset by designing a proper $\delta$. We will present in the following part a threat model of the paper.
Threat model.  The producer of UEs (i.e., defender) has white-box access to a surrogate classifier $f_{\hat{\theta}}$ and a protectee dataset ${\mathbb{D}}_{s}\sim{\mathcal{D}}$. Additionally, either a clean test dataset ${\mathbb{D}}$ or the domain ${\mathcal{D}}$ of ${\mathbb{D}}_{s}$ is revealed to the defender. Note that it is not necessary for the defender to access ${\mathbb{D}}$ if it is privately held by a third party. Instead, the defender can sample a dataset $\hat{{\mathbb{D}}}$ from ${\mathcal{D}}$ to compute a certified generalization learnability which serves as a generalization upper bound on the certified learnability. The defender can alter the parameters of $f_{\hat{\theta}}$ and modify ${\mathbb{D}}_{s}$ at will. The defender crafts an unlearnable version ${\mathbb{D}}_{s}\oplus\delta$ of ${\mathbb{D}}_{s}$ and releases the unlearnable dataset rather than the unprotected one to the public. Besides, defenders can either train surrogate classifiers on generated ${\mathbb{D}}_{s}\oplus\delta$ or simultaneously with $\delta$ on ${\mathbb{D}}_{s}$. However, the defender has zero access to any potential unauthorized classifiers and is unaware of the training procedure for the unauthorized classifiers.
@@
Among these adversaries, GA encompasses almost all attackers in real-world scenarios. CA denotes the subset of attackers from GA that can be provably mitigated by the defense. SA serves as a tool to gauge the robustness of different UEs against GAs who cannot be provably invalidated, as well as to assess the tightness of our defense against CAs. The defender aims to expand $\hat{\Theta}$ to improve the proportion of CAs within GAs. We will introduce in Section VI a method for augmenting $\hat{\Theta}$.
IV Overview
We will provide a brief overview over the crux of learnability certification and PUE generation in this section. Our first desideratum is measuring how well an unlearnable dataset can sabotage those illegal but unknown classifiers by certifying a possibly best learning outcome on this dataset. In a nutshell, a surrogate classifier selected based on the unlearnable set ${\mathbb{D}}_{s}\oplus\delta$ can be viewed as a hypothesis of the data distribution. Intuitively, if the distribution of the unlearnable set is highly disjoint from that of clean data, the classifier would demonstrate a significant generalization error on the clean test dataset. In contrast, if the unlearnable set is suboptimal, the classifier might still be able to generalize to the clean dataset and thus induce a lower generalization error. Herein, we convert the certification problem to certifying a surrogate classifier appropriately modeling the distribution of UEs.
Second, we design a simple method for selecting the surrogate classifier used in the certification. Importantly, we emphasize that the two entities, the UE and the surrogate classifier, serve different objectives. On the side of UEs, their purpose is upfront — to minimize the learnability by lowering the clean testing accuracy of any classifiers trained on them. Instead, the surrogate should 1) resemble the distributional characteristics of UEs and 2) have a high chance of discovering better clean test accuracy when its parameters are perturbed. In lieu of directly using classifiers trained by UEs as surrogates, we design a better way to construct surrogates meeting these two criteria.
@@ V Learnability Certification
V-A Quantile Parametric Smoothing
@@
Definition 2 (Quantile Parametric Smoothing function).
@@
$h_{q}(\theta)=\inf\ \{t\ |\ \Pr_{\epsilon}[A_{{\mathbb{D}}}(\theta+\epsilon)\leq t]\geq q\},$ (7)
@@
The definition connects to percentile smoothing used in certified robustness [32, 33]. In lieu of randomizing the inputs of classifiers, QPS takes the parameters (i.e., weights) of classifiers as its variables and randomizes the parameters.
@@
V-B Certified Learnability
To apply the QPS function to obtain a learnability certification, we explicitly define $A_{{\mathbb{D}}}$ as a performance metric in this section. Without loss of generality, such $A_{{\mathbb{D}}}(\hat{\theta})$ can be the top-1 accuracy function of a classifier ${f}_{\hat{\theta}}$ mapping its input $x$ to a label $y$, where $(x,y)\in{\mathbb{D}}$. That is
$A_{{\mathbb{D}}}(\hat{\theta}):=\mathbb{E}_{(x,y)\sim{\mathbb{D}}}\mathds{1}[{f}_{\hat{\theta}}(x)=y].$ (8)
@@ V-B Certified Learnability
Figure 3: An illustration of the certified $(q,\eta)$-Learnability.
@@
Theorem 1 (Perturbation bound on QPS).
@@
${h}_{q}(\hat{\theta}+\upsilon)\leq\inf\ \{t\ |\ \Pr_{\epsilon}[A_{{\mathbb{D}}}(\hat{\theta}+\epsilon)\leq t]\geq\overline{q}\},\ \forall\ \|\upsilon\|\leq\eta,$ (9)
@@
We defer the proof of Theorem 1 to Appendix -B. The theorem states that the return of the QPS function is bounded with respect to the perturbation $\upsilon$ in its input. Recall that ${h}_{q}(\hat{\theta}+\upsilon)$ is actually an upper bound of the clean test accuracy for classifiers whose parameters $\theta\sim\mathcal{N}(\hat{\theta}+\upsilon,\sigma^{2}I)$, with probability at least $q$. Therefore, the right-hand side of Inequality 9 actually guarantees the best possible clean test accuracy for classifiers with parameters from the subspace $\hat{\Theta}:=\{\theta\ |\ \theta\sim\mathcal{N}(\hat{\theta}+\upsilon,\sigma^{2}I),\ \|\upsilon\|\leq\eta\}$. We name such $\hat{\Theta}$ as a certified parameter set. Though it seems that $\hat{\Theta}$ covers all parameters in $\mathbb{R}^{d}$, we only consider those $\theta$ having a sufficiently large probability of being sampled.
Based on Theorem 1, we have the formal definition of $(q,\eta)$-Learnability as follows.
Definition 3 ($(q,\eta)$-Learnability).
@@
$l_{(q,\eta)}(\hat{\Theta};{\mathbb{D}}_{s}\oplus\delta)=\inf\ \{t|\Pr_{\epsilon}[A_{{\mathbb{D}}}(\hat{\theta}+\epsilon)\leq t]\geq\overline{q}\},$ (10)
@@
It should be noted that, though ${\mathbb{D}}_{s}\oplus\delta$ does not explicitly appear in the function, $\hat{\theta}$ is selected by a learning function $\hat{\theta}=\Gamma({\mathbb{D}}_{s}\oplus\delta)$ given ${\mathbb{D}}_{s}\oplus\delta$. Henceforth, $\hat{\theta}$ can be viewed as a function of ${\mathbb{D}}_{s}\oplus\delta$ in this case. In practice, we can sample $A_{{\mathbb{D}}}(\hat{\theta}+\epsilon)$ using Monte Carlo and obtain the accuracy in the empirical $\overline{q}$-th quantile. Next, a Binomial confidence interval upper bound of the accuracy can be adopted as $l_{(q,\eta)}(\hat{\Theta};{\mathbb{D}}_{s}\oplus\delta)$ with a confidence level of $1-\alpha$. Detailed steps of computing the confidence interval upper bound are in Appendix -A.
@@
V-C Certification Algorithm
We summarize the certification algorithm in this section. We first train a surrogate classifier $f_{\hat{\theta}}$ of which the parameters $\hat{\theta}$ will be randomized for $n$ times over a Gaussian distribution to calculate the QPS function. For each randomization, the resulting classifier with parameters $\hat{\theta}+\epsilon$ is evaluated on the clean test set ${\mathbb{D}}$ such that $A_{{\mathbb{D}}}(\hat{\theta}+\epsilon)$ is obtained. Next, $\overline{q}$ can be theoretically computed based on Theorem 1. Finally, we compute the confidence interval of the $A_{{\mathbb{D}}}(\hat{\theta}+\epsilon)$ in the $\overline{q}$-th quantile to obtain the upper bound of the interval as $l_{(q,\eta)}(\hat{\Theta};{\mathbb{D}}_{s}\oplus\delta)$. The detailed certification process is depicted in Algorithm 1 and Algorithm 2. In addition, a meticulous description is in Appendix -A.
There are a few important things to be noticed. A higher $q$ is recommended for obtaining a learnability upper bound with high probability, and a larger $\eta$ helps establish a more inclusive certified parameter set. In practice, both parameters can be optimized through grid search. Moreover, the $\sigma$ of the parametric smoothing noise should be selected with caution. There is an intuition behind randomizing the surrogate weights and computing the QPS function — some of the randomized classifiers can outperform the surrogate classifier on the clean test set. Notwithstanding, we find that, when an improperly large $\sigma$ is employed, the number of randomized classifiers having a higher accuracy than that of the surrogate drops significantly. The possible reason for this phenomenon could be that the shortcuts added to the UEs cannot fully conceal the mapping information between input samples and their labels. Hence, the surrogate trained on the unlearnable dataset, though not adequately, is somehow fit to the clean data distribution as well. Large noises added to the parameters thus not only shift the surrogate away from the distribution of UEs but also move it off the clean data manifold. This assumption is further verified by the recovery attack, which discovers classifiers with satisfactory performance on clean data in the vicinity of the poisoned classifier. Therefore, we will further introduce the methods for constructing sound surrogates for tighter certification.
func QUpperBound
@@ V-C Certification Algorithm
$\overline{q}\leftarrow\Phi(\Phi^{-1}(q)+\frac{\eta}{\sigma})$
$\underline{k},\ \overline{k}\ \leftarrow\ \lceil n*\overline{q}\rceil,\ n$
@@
for $k\in\{\underline{k},\underline{k}+1,...,\overline{k}\}$ do
       if $\textsc{Binomial}(n,k,\overline{q})>1-\alpha$ then
             $k^{*}\ \leftarrow k$
      else
             $Continue$
if $k^{*}\neq 0$ then
@@
else
       Abstain
func Binomial
@@
$\textsc{Conf}\ \leftarrow\ \sum_{i=1}^{k}\binom{n}{k}(\overline{q})^{i}(1-\overline{q})^{n-i}$
@@
Initialize $a$
for $i\in 1,...,n$ do
       $\theta\ \leftarrow\ \hat{\theta}+\epsilon$, $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$
       Evaluate $A_{{\mathbb{D}}}({\theta})$ on ${\mathbb{D}}$
       Append $A_{{\mathbb{D}}}({\theta})$ to $a$
$a\ \leftarrow\ Sort(a)$
$k\ \leftarrow\ \textsc{QUpperBound}(n,\alpha,\sigma,\eta,q)$
$t\ \leftarrow\ a_{k}$
@@ V-D Generalization of the Certification
As a solution, the defender can sample a test set $\hat{{\mathbb{D}}}$ from the accessible domain ${\mathcal{D}}$ and certify a generalization $(q,\eta)$-Learnability which bounds the learnability on test sets drawn from ${\mathcal{D}}$. By Hoeffding’s Inequality, $A_{{\mathcal{D}}}(\hat{\theta})\leq A_{\hat{{\mathbb{D}}}}(\hat{\theta})+\sqrt{\frac{1}{2N}\log(\frac{2}{\beta})}$ exists with probability at least $1-\beta$. The right-hand side of the inequality can be used in the QPS function instead of $A_{{\mathbb{D}}}(\hat{\theta})$. Combining with the union bound, a certified generalization $(q,\eta)$-Learnability can be acquired as $l_{(q,\eta)}(\hat{\Theta};{\mathbb{D}}_{s}\oplus\delta)+\sqrt{\frac{1}{2N}\log(\frac{2n}{\beta})}$, where $n$ is the number of noise draws. However, Hoeffding’s Inequality does not account for factors that improve generalization learnability. Therefore, we explore alternative approaches to establish the relationship between generalization learnability and the key components involved in the certification process. Notably, classifiers sampled around a surrogate relate to PAC-Bayesian predictors (e.g., Gibbs classifiers) whose parameters are randomly drawn from a distribution [34]. We can thus bound the expected generalization accuracy under random weight perturbation through the following corollary.
Corollary 1 (Expected generalization accuracy under parametric smoothing noise).
@@
$\mathbb{E}_{\epsilon}[A_{{\mathcal{D}}}(\hat{\theta}+\epsilon)]\geq\mathbb{E}_{\epsilon}[{A}_{\hat{{\mathbb{D}}}}(\hat{\theta}+\epsilon)]-\sqrt{\frac{\frac{\|\hat{\theta}\|^{2}}{\sigma^{2}}+\ln\frac{N}{\alpha}}{2(N-1)}},$ (12)
@@
The corollary indicates that the expected generalization accuracy over the clean data distribution is lower bounded by the expected empirical accuracy minus a function of $\hat{\theta}$, $\sigma$ and $N$. Although the corollary bounds the expectation rather than the $q$-th quantile, it connotes that a $f_{\hat{\theta}}$ constantly producing correct predictions on $\hat{{\mathbb{D}}}$ also has a high generalization accuracy, which evinces the inferiority of the UEs resulting $\hat{\theta}$. Furthermore, if the expectation of ${A}_{\hat{{\mathbb{D}}}}(\hat{\theta}+\epsilon)$ is robust to varying $\sigma$, the expected generalization accuracy would grow with increasing $\sigma$. For appropriately skewed accuracy distributions, this result implies that an $f_{\hat{\theta}}$ which endures large parametric noise can certify $(q,\eta)$-Learnability that generalizes better to ${\mathcal{D}}$.
@@
[backgroundcolor=grey!10,rightline=true,leftline=true,topline=true,bottomline=true,roundcorner=1mm,everyline=false,nobreak=false] Remark.  We defined the QPS function of accuracy scores obtained by a set of classifiers randomized from a surrogate classifier fit to UEs ${\mathbb{D}}_{s}\oplus\delta$. The QPS function returns the value in the $q$-th quantile of all accuracy scores. The output of the QPS function can be further upper bounded by Theorem 1. The upper bound of the accuracy in the $\overline{q}$-th quantile is treated as the $(q,\eta)$-Learnability of ${\mathbb{D}}_{s}\oplus\delta$. The key message of the certified $(q,\eta)$-Learnability is that — if an unauthorized classifier has its trained parameters in the certified parameter set $\hat{\Theta}:=\{\theta\ |\ \theta\sim\mathcal{N}(\hat{\theta}+\upsilon,\sigma^{2}I),\ \|\upsilon\|\leq\eta\}$, then its clean test accuracy will not exceed the $(q,\eta)$-Learnability, with probability at least $q$. Moreover, if ${\mathbb{D}}$ is hidden from the defender, a generalization $(q,\eta)$-Learnability can be certified by sampling a test set from ${\mathcal{D}}$ by the defender.
VI Provably Unlearnable Examples